# Design Patterns for the Common Good: Building Better Technologies Using the Wisdom of Virtue Ethics

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXsubmissionid: 4471ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
  • Seven design patterns inspired by Catholic Social Teaching to build social media more in accordance with human flourishing, articulating a vision of what such ethically designed online spaces would look like.
  • The first consideration of how Catholic values can influence design, contributing to the growing field of faith and spirituality within HCI and bringing a new ethical dimension into discussions of technology design.
@@ 2.2. Philosophy and Spirituality in Technology Design
Other work has imagined technology design inspired by religious traditions. A recent ACM Interactions issue dedicated to spirituality in design included articles on how Judaism (Hammer and Reig, 2022), Christianity (Hiniker and Wobbrock, 2022), and religiosity at large (Toyama, 2022; Naqshbandi et al., 2022), could inform technology design. For example, Hiniker and Wobbrock (Hiniker and Wobbrock, 2022) proposed that technology designed with Christian principles in mind would promote relationship with God, others, and the environment, and would be at odds with the attention economy — the primary economic model for many of our technologies today. Similarly, Hammer and Reig discussed how Jewish conceptions of obligations can provide a novel conceptual approach and overarching design implications for online speech (Hammer and Reig, 2022). Additionally, Toyama discussed how designing technologies to foster the inner attention valued by religious traditions could “counter technological excess and corruption” (Toyama, 2022). Finally, Naqshbandi et al. argued that capturing faith, religion, and spirituality-based motives is important for advancing design justice and getting a more complete picture of human motives and conduct (Naqshbandi et al., 2022). Beyond the Interactions issue, other HCI work has demonstrated that self-tracking apps built with Buddhist principles enable users to develop better emotional and ethical awareness as well as less technology addiction compared to standard self-tracking apps (McGuire, 2020).
All of the works mentioned argue that philosophical or religious traditions should inform technology design. However, the proposed design approaches remain at theoretical suggestions. To our knowledge, our work is the first to translate principles from philosophical or religious traditions into practical design recommendations.
@@ 2.4. Design Patterns
Design patterns for social media exist (Crumlish and Malone, 2015) (Hussein et al., 2011). However, while these design patterns capture the key building blocks for how to build easy-to-use social media applications, these patterns do not make any claims about the ethics or digital well-being they promote. (One study of designs to decrease screen time on Twitter recommends these suggestions to promote digital well-being be cataloged as design patterns, but to our knowledge they have not been (Zhang et al., 2022).) We believe this work is the first to document design patterns for social media that have a particular ethical backing.
@@ 3.1. Virtue Ethics
The philosopher Shannon Vallor popularized considering the ethics of emerging technologies in terms of virtue ethics through her 2016 book Technology and the Virtues (Vallor, 2018). In this book, Vallor analyzes the virtues from the Aristotelian, Buddhist, and Confucianist virtue traditions to identify which are important for living well with technology today. Vallor then considers the ethics of technologies including social media, surveillance, robots, and human enhancement in light of these virtues. Inspired by Vallor, a small number of other scholars have considered how virtue ethics can inform technology design; for example, using virtue ethics to consider the intellectual virtues required for virtuous online search, and proposing high-level design implications from these virtues (Gorichanaz, 2024). To our knowledge, this paper is the first to recommend designs that resonate with the virtue ethics tradition at the level of specificity of design patterns.
@@ 3.2. Catholic Social Teaching
As a case study of our proposed approach we will identify design patterns that resonate with the Catholic tradition. Specifically, the virtues we will consider in our design pattern identification are the main principles of Catholic Social Teaching: the area of Catholic doctrine concerned with upholding human dignity and the common good in society.
@@ 4. Approach: Virtue-Guided Technology Design
Conceptual Inquiry: This stage entails considering how these virtues may play out in the context of the technology one desires to design. According to Friedman et al. (Friedman et al., 2013), “Conceptual investigations do not by themselves involve costly empirical analyses, but instead thoughtful consideration of how stakeholders might be socially impacted by one’s technological designs.” In our example, the designer may conclude after thoughtful consideration that temperance in video streaming would not encourage users to watch content for hours on end.
Technical Inquiry: After working definitions for how the chosen values play out in an online context are established, the process moves on to the Technical Inquiry. First, identify already-existing technologies that are similar to the technology the design patterns are being cataloged for. Then, analyze the technical features of these example technologies, considering how they hold up in light of the conceptual definitions of the identified virtues. This results in a list of features that either promote or hinder the chosen virtues. In our example, the designer might consider Netflix, and note that its autoplay feature violates temperance in the way identified in their conceptual inquiry.
@@
As a proof of concept of our proposed approach, we use Virtue-Guided Technology Design to create design patterns for social media platforms inspired by the virtues of Catholic Social Teaching.
@@ 5.2. Technical Inquiry: Which Social Media Design Features Uphold or Violate Catholic Social Teaching
In the technical inquiry, we identified the designs used in various social communications platforms and checked them against the conceptual definitions of Catholic Social Teaching principles to identify patterns of virtuous (or vicious) designs. We analyzed the following social communications platforms in light of the conceptual values: X (formerly known as Twitter), TikTok, BeReal, Gmail, and Discord. We selected these platforms because while they are all social technologies, they encapsulate a wide variety of features and mediums through which communication occurs. First, for each platform we listed out a comprehensive set of features from that platform. Then, for each feature we considered how it holds up in terms of the conceptual definitions of each theme of Catholic Social Teaching as articulated in Virtue in Virtual Spaces. Finally, we categorized features by the principles of Catholic Social Teaching they upheld or violated.
[…]
We then documented design patterns based on how we observed patterns of features upholding or violating the principles of Catholic Social Teaching according to their conceptual definitions. An illustration of this process for one design pattern can be seen in Figure 4; illustrations for the other patterns can be found in this paper’s Supplemental Material. After generating eleven potential patterns, we had iterative group discussions to choose a subset of patterns that best embodied the principles of Catholic Social Teaching, emphasized interpersonal communication, and/or had the potential to generate thoughtful discussions. A preliminary set of seven patterns were chosen, including: 1) Chats Over Feeds, 2) Friends Over Followers, 3) Moderated Mingling, 4) Clear Algorithmic Comprehension, 5) Moderated Entry, 6) No Lurking, and 7) Notification Intentionality. Additional information about each pattern can be found in Figure 5.
@@ 6.1.2. Mapping the Participant-Identified Concerns to Violations of Catholic Social Teaching
Challenges of moderation This concern encompasses difficulties of rule-setting in moderation and concerns that moderators may not be benevolent. P4 raised the concern, “what if people are trying to contribute to the conversation, but the community guidelines decide that’s off topic or that’s hate speech?” P15 additionally raised a concern about “abuses of power” from moderators. While these concerns are often mitigated by good moderators, the larger the group the more inevitable these issues become (Gillespie, 2018). Life and dignity of the human person pertains to the ability to have authentic dialogue, and option for the poor and vulnerable pertains to user well-being, both of which could be violated if moderators are not benevolent or are not able to enforce good rules (either because of their own incompetence or because the group exists at too large a scale). Challenges of moderation was mentioned for the following patterns: Moderated Mingling (N=14), and Moderated Entry (N=4).
@@
We gave the design patterns violation scores for each principle of Catholic Social Teaching by adding together the number of times each of the concerns violating that particular principle was mentioned for that pattern. For example, Chats Over Feeds had six participants mention increased moral distance and 15 participants mention limited learning, for a total life and dignity violation score of 21. Because implementation concerns, challenges of moderation, and undesirable self-knowledge did not correspond to violations of the principles of Catholic Social Teaching, we did not consider these concerns in the scores. The violation scores will be used in Section 6.1.3 along with the embodiment scores calculated in Section 6.1.1 to compute a net embodiment score for each pattern.
@@ 6.1.3. Overall, do the patterns embody or violate Catholic Social Teaching?
No Lurking was intended to embody the call to family, community, and participation. The heat map indicates it additionally embodies life and dignity of the human person and option for the poor and vulnerable.
@@ 6.2. Do people like the patterns? Would they use a social media designed with these patterns?
We determined the answer to RQ3 by considering how many participants rated each pattern as “net positive,” “net negative,” or “net neutral.” In the interviews “net neutral” was primarily interpreted as “ambivalent” by the participants; in other words, the pattern had both strong values and strong concerns that the participant could not choose between positive and negative. Because of this, we will refer to “net neutral” responses as “ambivalent” in our results.
@@ 6.3. Do tech professionals think these patterns have a place in industry?
One a more hopeful note, two participants expressed that while big tech may not adopt these patterns, smaller socially-conscious companies or nonprofits would value them. In fact, P13, who was a product designer for a Catholic company building a chat platform, noted that our design patterns resonated with the company’s design priorities:
@@
Two of the eight participants who were asked about our proposed patterns’ place in industry were atheist/agnostic. These two participants were additionally asked about how the Catholic Social Teaching themes inspiring the patterns resonated with their personal value systems. Both participants noted that the themes of Catholic Social Teaching resonated with their personal values despite not being Christian. One participant noted that from their knowledge of Catholicism, some of the specific details of how a Catholic may live out the Catholic Social Teaching themes may be different from how they would; for example, for a Catholic, the call to family, community, and participation may have a stronger emphasis on the nuclear family, whereas for this participant, they may put less of an emphasis on blood or legal relation. However, the over-arching themes still resonated.
@@ 7.2. Adoption of the Patterns
Although these patterns were identified to be in accord with Catholic virtues, we believe they would be adopted by non-Catholics as well. The atheist and agnostic study participants expressed that the virtues of Catholic Social Teaching underpinning these design patterns resonated with their personal value systems, which indicates that Catholic Social Teaching in particular could be easily brought into multicultural settings. This is fitting given that Catholic Social Teaching is not meant just for Catholics, but is intended to bring all people concerned with social justice into conversation. The design patterns proposed in this paper embody the principles of Catholic Social Teaching at a general enough level such that they resonated cross-culturally amongst our study participants, but tensions could arise if future patterns were created in a more specifically Catholic sense. To use the example provided by one of our participants, if a design pattern that upholds the call to family, community, and participation over-emphasizes connecting with one’s nuclear family, this pattern may not resonate with communities that focus less on the importance of blood or legal relation. When employing Virtue-Guided Technology Design, we encourage the consideration of the wider resonance of the chosen traditions’ virtues across other value systems.
@@ 7.3. The Vision Put Forth By These Patterns
The vision put forth by our seven design patterns is a vision of intentional conversations, either in unmoderated small groups or in moderated large-group chat rooms. While participants generally liked this vision, as exhibited by many “net positive” responses to the patterns, they did not want this type of interaction to be the only type available. This is seen in the high number of ambivalent responses, especially for the Chats Over Feeds and Friends Over Followers patterns, and also through the concerns expressed. The concern mentioned the most across all patterns was limited learning, which encompasses not being able to see other viewpoints or gather information. The limited learning concern violated the principle of solidarity, which we found was the primary theme of Catholic Social Teaching that our patterns violated. Thus we see a dichotomy emerge between fostering intentional conversation and gathering information or seeing new ideas. How do we reconcile the fact that smaller groups and more guardrails can help to foster more intentional conversations, but these very same guardrails can curtail the spread of ideas? The Moderated Mingling pattern was intended to foster this large group sharing of ideas. However, it seems that participants did not find this to be a fully satisfying way of sharing ideas widely and they expressed doubts that moderation would foster a truly free sharing of ideas.
@@ 7.4. Iterating on the Design Patterns for Further Value Discovery
This is especially important given that the translation of virtues to technical features is a subjective process. In this work, the authors relied on our own previous experience and engagement with Catholic Social Teaching when identifying its virtues in designs. The presence of bias from the researchers in their interpretation of values does not render the work invalid: we have to start somewhere. Rather, this is an invitation for greater conversation surrounding the translation of these virtues into technology designs. This is especially important given the cultural diversity within Catholicism itself. Although the Catholic Church holds certain doctrines to be true, the way these doctrines are interpreted and lived out may vary across cultures. For example, Catholicism in the United States may look different from Latin American Catholicism, which may look different from African Catholicism. The design patterns put forth in this paper are merely a starting point based on our interpretations. In addition to our hope that researchers will use our Virtue-Guided Technology Design process with other faith traditions, we also hope that other researchers from around the world will use this process to further engage with Catholic Social Teaching, helping to draw out more of its richness because of their unique experiences and perspectives.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713546isbn: 979-8-4007-1394-1/25/04submissionid: 4471ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
  • Seven design patterns inspired by Catholic Social Teaching to build social media more in accordance with human flourishing, articulating one vision of what such ethically designed online spaces could look like.
  • The first consideration of how Catholic values can influence design, contributing to the growing field of faith and spirituality within HCI and bringing a new ethical perspective into discussions of technology design.
@@ 2.2. Philosophy and Spirituality in Technology Design
Other work has imagined technology design inspired by religious traditions. A recent ACM Interactions issue dedicated to spirituality in design included articles on how Judaism (Hammer and Reig, 2022), Christianity (Hiniker and Wobbrock, 2022), and religiosity at large (Toyama, 2022; Naqshbandi et al., 2022) could inform technology design. For example, Hiniker and Wobbrock (Hiniker and Wobbrock, 2022) proposed that technology designed with Christian principles in mind would promote relationship with God, others, and the environment, and would be at odds with the attention economy — the primary economic model for many of our technologies today. Similarly, Hammer and Reig discussed how Jewish conceptions of obligations can provide a novel conceptual approach and overarching design implications for online speech (Hammer and Reig, 2022). Additionally, Toyama discussed how designing technologies to foster the inner attention valued by religious traditions could “counter technological excess and corruption” (Toyama, 2022). Finally, Naqshbandi et al. argued that capturing faith, religion, and spirituality-based motives is important for advancing design justice and getting a more complete picture of human motives and conduct (Naqshbandi et al., 2022). Beyond the Interactions issue, other HCI work has demonstrated that self-tracking apps built with Buddhist principles enable users to develop better emotional and ethical awareness as well as resist technology addiction compared to standard self-tracking apps (McGuire, 2020).
All of the works mentioned argue that philosophical or religious traditions should inform technology design. However, most of the proposed design approaches remain at theoretical suggestions. In contrast, our approach allows for principles from diverse ethical traditions to be translated into practical design recommendations.
@@ 2.4. Design Patterns
Design patterns for social media exist (Crumlish and Malone, 2015) (Hussein et al., 2011). However, while these design patterns capture the key building blocks for how to build easy-to-use social media applications, these patterns do not make any claims about the ethics or digital well-being they promote. (One study of designs to decrease screen time on Twitter recommends these suggestions to promote digital well-being be cataloged as design patterns, but to our knowledge they have not been (Zhang et al., 2022).) In contrast, this work documents design patterns for social media that have a particular ethical backing.
@@ 3.1. Virtue Ethics
The philosopher Shannon Vallor popularized considering the ethics of emerging technologies in terms of virtue ethics through her 2016 book Technology and the Virtues (Vallor, 2018). In this book, Vallor analyzes the virtues from the Aristotelian, Buddhist, and Confucianist virtue traditions to identify which are important for living well with technology today. Vallor then considers the ethics of technologies including social media, surveillance, robots, and human enhancement in light of these virtues. Inspired by Vallor, a small number of other scholars have considered how virtue ethics can inform technology design; for example, using virtue ethics to consider the intellectual virtues required for virtuous online search, and proposing high-level design implications from these virtues (Gorichanaz, 2024). To our knowledge, this paper is the first to recommend designs that align with the virtue ethics tradition at the level of specificity of design patterns.
@@ 3.2. Catholic Social Teaching
As a case study of our proposed approach we will identify design patterns that align with the Catholic tradition. Specifically, the virtues we will consider in our design pattern identification are the main principles of Catholic Social Teaching: the area of Catholic doctrine concerned with upholding human dignity and the common good in society.
@@ 4. Approach: Virtue-Guided Technology Design
Conceptual Inquiry: This stage entails considering how these virtues may play out in the context of the technology one desires to design. According to Friedman et al. (Friedman et al., 2013), “Conceptual investigations do not by themselves involve costly empirical analyses, but instead thoughtful consideration of how stakeholders might be socially impacted by one’s technological designs.” In our example, the designer may conclude after thoughtful consideration that a video streaming platform embodying temperance would not encourage users to watch content for hours on end.
Technical Inquiry: After working definitions for how the chosen values play out in an online context are established, the process moves on to the Technical Inquiry. First, identify already-existing technologies that are similar to the technology the design patterns are being cataloged for. Then, analyze the technical features of these example technologies, considering how they hold up in light of the conceptual definitions of the identified virtues. This results in a list of features that either promote or hinder the chosen virtues. In our example, the designer might consider Netflix, and note that its autoplay feature violates temperance in the way she identified in the conceptual inquiry.
@@
[…]
As a proof of concept of our proposed approach, we use Virtue-Guided Technology Design to identify design patterns for social media platforms inspired by the virtues of Catholic Social Teaching.
@@ 5.2. Technical Inquiry: Which Social Media Design Features Uphold or Violate Catholic Social Teaching
In the technical inquiry, we identified the designs used in various social communications platforms and checked them against the conceptual definitions of Catholic Social Teaching principles to identify patterns of virtuous (or unvirtuous) designs. We analyzed the following social communications platforms in light of the conceptual definitions: X (formerly known as Twitter), TikTok, BeReal, Gmail, and Discord. We selected these platforms because while they are all social technologies, they encapsulate a wide variety of features and mediums through which communication occurs. First, for each platform we listed out a comprehensive set of features from that platform. Then, for each feature we considered how it holds up in terms of the conceptual definitions of each principle of Catholic Social Teaching as articulated in Virtue in Virtual Spaces. Finally, we categorized features by the principles of Catholic Social Teaching they upheld or violated.
We then documented design patterns based on how we observed patterns of features upholding or violating the principles of Catholic Social Teaching according to their conceptual definitions. An illustration of this process for one design pattern can be seen in Figure 4; illustrations for the other patterns can be found in the appendix. After generating eleven potential patterns, we had iterative group discussions to choose a subset of patterns that best embodied the principles of Catholic Social Teaching, emphasized interpersonal communication, and/or had the potential to generate thoughtful discussions. A preliminary set of seven patterns were chosen, including: 1) Chats Over Feeds, 2) Friends Over Followers, 3) Moderated Mingling, 4) Clear Algorithmic Comprehension, 5) Moderated Entry, 6) No Lurking, and 7) Notification Intentionality. Additional information about each pattern can be found in Figure 5.
@@ 6.1.2. Mapping the Participant-Identified Concerns to Violations of Catholic Social Teaching
Challenges of moderation (life and dignity of the human person, option for the poor and vulnerable): This concern encompasses difficulties of rule-setting in moderation and concerns that moderators may not be benevolent. P4 raised the concern, “what if people are trying to contribute to the conversation, but the community guidelines decide that’s off topic or that’s hate speech?” P15 additionally raised a concern about “abuses of power” from moderators. While these concerns are often mitigated by good moderators, the larger the group the more inevitable these issues become (Gillespie, 2018). Life and dignity of the human person pertains to the ability to have authentic dialogue, and option for the poor and vulnerable pertains to user well-being, both of which could be violated if moderators are not benevolent or are not able to enforce good rules (either because of their own incompetence or because the group exists at too large a scale). Challenges of moderation was mentioned for the following patterns: Moderated Mingling (N=14), and Moderated Entry (N=4).
@@
We gave the design patterns violation scores for each principle of Catholic Social Teaching by adding together the number of times each of the concerns violating that particular principle was mentioned for that pattern. For example, Chats Over Feeds had six participants mention increased moral distance and 15 participants mention limited learning, for a total life and dignity violation score of 21. Because implementation concerns and undesirable self-knowledge did not correspond to violations of the principles of Catholic Social Teaching, we did not consider these concerns in the scores. The violation scores will be used in Section 6.1.3 along with the embodiment scores calculated in Section 6.1.1 to compute a net embodiment score for each pattern.
@@ 6.1.3. Overall, do the patterns embody or violate Catholic Social Teaching?
No Lurking was intended to embody the call to family, community, and participation. The heat map indicates it additionally embodied life and dignity of the human person and option for the poor and vulnerable.
[…]
@@ 6.2. Do people like the patterns? Would they use a social media designed with these patterns?
We determined the answer to RQ3 by considering how many participants rated each pattern as “net positive,” “net negative,” or “net neutral.” In the interviews “net neutral” was primarily interpreted as “ambivalent” by the participants; in other words, the pattern had both strong values and strong concerns such that the participant could not choose between positive and negative. Because of this, we will refer to “net neutral” responses as “ambivalent” in our results.
@@ 6.3. Do tech professionals think these patterns have a place in industry?
On a more hopeful note, two participants expressed that while big tech may not adopt these patterns, smaller socially-conscious companies or nonprofits would value them. In fact, P13, who was a product designer for a Catholic company building a chat platform, noted that our design patterns resonated with the company’s design priorities:
@@
Two of the eight participants who were asked about our proposed patterns’ place in industry were atheist/agnostic. These two participants were additionally asked about how the Catholic Social Teaching principles inspiring the patterns resonated with their personal value systems. Both participants noted that the principles of Catholic Social Teaching resonated with their personal values despite not being Christian. One participant noted that from their knowledge of Catholicism, some of the specific details of how a Catholic may live out the Catholic Social Teaching principles may be different from how they would; for example, for a Catholic, the call to family, community, and participation may have a stronger emphasis on the nuclear family, whereas for this participant, they may put less of an emphasis on blood or legal relation. However, the over-arching ideas still resonated.
@@ 7.2. Adoption of the Patterns
Although these patterns were identified to be in accord with Catholic virtues, we believe they would be adopted by non-Catholics as well. The atheist and agnostic study participants expressed that the principles of Catholic Social Teaching underpinning these design patterns resonated with their personal value systems, which indicates that Catholic Social Teaching in particular could be easily brought into multicultural settings. This is fitting given that Catholic Social Teaching is not meant just for Catholics, but is intended to bring all people concerned with social justice into conversation. The design patterns proposed in this paper embody the principles of Catholic Social Teaching at a general enough level such that they resonated cross-culturally amongst our study participants, but tensions could arise if future patterns were created in a more specifically Catholic sense. To use the example provided by one of our participants, if a design pattern that upholds the call to family, community, and participation over-emphasizes connecting with one’s nuclear family, this pattern may not resonate with communities that focus less on the importance of blood or legal relation. When employing Virtue-Guided Technology Design, we encourage the consideration of the wider resonance of the chosen traditions’ virtues across other value systems.
@@ 7.3. The Vision Put Forth By These Patterns
The vision put forth by our seven design patterns is a vision of intentional conversations, either in unmoderated small groups or in moderated large-group chat rooms. While participants generally liked this vision, as exhibited by many “net positive” responses to the patterns, they did not want this type of interaction to be the only type available. This is seen in the high number of ambivalent responses, especially for the Chats Over Feeds and Friends Over Followers patterns, and also through the concerns expressed. The concern mentioned the most across all patterns was limited learning, which encompasses not being able to see other viewpoints or gather information. The limited learning concern violated the principle of solidarity, which we found was the primary principle of Catholic Social Teaching that our patterns violated. Thus we see a dichotomy emerge between fostering intentional conversation and gathering information or seeing new ideas. How do we reconcile the fact that smaller groups and more guardrails can help to foster more intentional conversations, but these very same guardrails can curtail the spread of ideas? The Moderated Mingling pattern was intended to foster this large group sharing of ideas. However, it seems that participants did not find this to be a fully satisfying way of sharing ideas widely and they expressed doubts that moderation would foster a truly free sharing of ideas.
@@ 7.4. Iterating on the Design Patterns for Further Value Discovery
This is especially important given that the translation of virtues to technical features is a subjective process. In this work, the authors relied on our own previous experience and engagement with Catholic Social Teaching when identifying its virtues in designs. The presence of bias from the researchers in their interpretation of the virtues does not render the work invalid: we have to start somewhere. Rather, this is an invitation for greater conversation surrounding the translation of these virtues into technology designs. This is especially important given the cultural diversity within Catholicism itself. Although the Catholic Church holds certain doctrines to be true, the way these doctrines are interpreted and lived out may vary across cultures. For example, Catholicism in the United States may look different from Latin American Catholicism, which may look different from African Catholicism. The design patterns put forth in this paper are merely a starting point based on our interpretations. In addition to our hope that researchers will use our Virtue-Guided Technology Design process with other faith traditions, we also hope that other researchers from around the world will use this process to further engage with Catholic Social Teaching, helping to draw out more of its richness because of their unique experiences and perspectives.